OpenStudy (anonymous):

(X-4)^2+4(X-4)-12=0 X^2+4X-12 (X+6)(X-2) X=-6,2 (X-4)-6=0 (X-4)+2=0 X-4=6...X=10? X-4=-2....X=2

OpenStudy (psymon):

Are we just checking your answer?

OpenStudy (anonymous):

\[X^ \frac{ 1 }{ 2 } + \sqrt{3X}^\frac{ 1 }{ 4 }-18\]

OpenStudy (psymon):

This is a little tricky, but it is still a quadratic and still is in the form of ax^2 + bx + c. If you notice, the power on the leading x is the same x^(1/4) squared. If that makes any sense. You see how?

OpenStudy (anonymous):

YES , IM STUCK ON SQRT3X BEING FACTORED WITH THE 18

OpenStudy (psymon):

Just by chance, you don't have a multiple choice thing, right? Just curious if the answer is supposed tobe funky.

OpenStudy (anonymous):

NOPE

OpenStudy (psymon):

Lots of caps o.o

OpenStudy (anonymous):

sorry

OpenStudy (psymon):

Well, because it's a square root 3, it looks like this is quadratic formula for sure. It's just being able to realize that it is a quadratic first.

OpenStudy (anonymous):

it is a Quadratic

OpenStudy (psymon):

I know.....sorry, didn't mean anything by saying that....either way, quadratic formula:

OpenStudy (psymon):

\[\frac{ -(3^{\frac{ 1 }{ 4 }})\pm \sqrt{(3^{\frac{ 1 }{ 4 }})^{2}-4(-18)} }{ 2 }\] Now it's just simplifying that mess as much as possible.